Five times the sum of a number and 3 is the same as 3 multiplied by 1 less than twice they number. What is t he number?: Five times the sum of a number and 3 is the same as 3 multiplied by 1 less than twice they number. What is t he number?
Answer by checkley77(3407)

5(X+3)=3(2X-1)
5X+15=6X-3
5X-6X=-3-15
-X=-18
X=-18/-1
X=18 ANSWER.
PROOF:
5(18+3)=3(2*18-1)
5*21=3(36-1)
105=3*35
105=105